# Dissecting Buchdahl’s limit: A surgeon’s guide to compact objects

### v1

@@ IV AdS Stars and Einstein Static stars
AdS stars are constituted by an anti-de Sitter interior glued through a thin shell to a Schwarzschild exterior. The geometry can be built through a cut and paste procedure with the help of Israel junction conditions. The interesting thing here is that the qualitative behaviour of this ad hoc geometry constructed through this surgery procedure appears naturally within the set of exact regular bilayered models, the only difference being that in that case the shell-like behaviour is absent and the transition between the inside core and the outside is smooth.
@@
Let us specify the construction in detail: Take a spherical shell with radius $R$, such that its proper area is $4\pi R^{2}$. Let the spacetime outside the shell be Schwarzschild with a mass $M$, and AdS inside the shell. We denote the Schwarzschild (outside) patch with a $+$ sign, and the AdS (inside) patch with a -- sign. The metric reads:
$ds^{2}_{+}=-f_{+}(r_{+})dt_{+}^{2}+\frac{dr_{+}^{2}}{f_{+}(r_{+})}+r_{+}^{2}d\Omega_{2}^{2},$ (42)
$ds^{2}_{-}=-f_{-}(r_{-})dt_{-}^{2}+\frac{dr_{-}^{2}}{f_{-}(r_{-})}+r_{-}^{2}d\Omega_{2}^{2},$ (43)
with $f_{+}(r_{+})=1-2M/r_{+}$ and $f_{-}(r_{-})=1+\absolutevalue{\lambda}r^{2}$, where $\lambda$ is a real parameter. We locate the shell at $r_{+}=r_{-}=R$. The matching is worked out in detail in Appendix B, and we find that the shell displays a distributional stress-energy tensor with the following densities and pressures:
@@
These models could serve as simple templates of black hole mimickers inspired by semiclassical effects. One central element responsible for both the features of their ringdown Cardoso and Pani (2019) and their associated shadows Ayzenberg et al. (2023) are the redshift and time delay effects suffered by light rays crossing their interiors. While the AdS star has a redshift function that decreases towards the centre, the one associated to the Einstein static star is fully constant. This would result in radically different crossing times for null rays and, in turn, to very different observational signatures both in the electromagnetic and gravitational-wave spectra.
@@ V Anisotropy
To isolate the effect of anisotropy in bypassing the Buchdahl limit, it is convenient to focus on a situation in which there are no violations of energy conditions arising from an interior core. For that purpose, we can analyze a toy model of a shell matching an exterior Schwarzschild spacetime with mass $M$ to an interior flat spacetime (i.e., Schwarzschild with $M=0$). The setup is the same as the one depicted in Fig. 7, but considering the core to be empty instead of AdS. This simply corresponds to setting the AdS constant $\lambda=0$ from previous results, and we find the following energy and tangential pressure densities:
@@
Given that the two of them have the same distributional behaviour, for our particular example this would imply
@@
We can compare these bounds with the shell model that we are considering, and we realize that the shell model begins to violate the DEC whenever the compactness is greater than $0.962M/R=$ 0.96, which still obeys the bound $2M/R\leq 0.974$ found by Ivanov Ivanov (2002). Further results that put bounds on the compactness of the objects with similar assumptions are the results reported in Barraco et al. (2003); Boehmer and Harko (2006); Lindblom (1984); Urbano and Veermäe (2019). In Barraco et al. (2003), under roughly speaking the same assumptions as Ivanov’s, it is shown that anisotropic models fall into two classes according to whether $p_{r}>p_{t}$ or $p_{t}>p_{r}$. In the case $p_{r}>p_{t}$, they manage to show that the $p_{r}$ function is strictly greater than the one in a fiduciary isotropic model with the same function $m(r)$. This means that these stars cannot be super-Buchdahl; actually, isotropic stars obeying DEC need to be less compact than Buchdahl Barraco and Hamity (2002). Regarding the case $p_{r}<p_{t}$, they show that the radial pressure is always bigger than the one in the fiduciary isotropic model with the same Misner-Sharp mass profile. Imposing the DEC, they derive bounds that are consistent with Ivanov’s. In Boehmer and Harko (2006), under some additional assumptions (about surface density behaviour motivated on physical grounds), slightly sharper bounds are proved. Finally, in Lindblom (1984); Urbano and Veermäe (2019) it is shown that imposing a causality condition, namely that the speed of sound is smaller than the speed of light, there is a maximum compactness that can be reached, which turns out to be below Buchdahl’s limit, namely $2M/R\lesssim 0.768$.
In any case, the simple example exhibited in this section shows that in the absence of energy conditions, anisotropic configurations can be made as compact as desired. Energy conditions set a limit to the compactness, but it is nevertheless far from Buchdahl’s, leaving the door open to having extremely compact objects that are close to forming event horizons, and are even allowed to display photon rings (since for that it is only required that $2M/R>2/3\sim 0.666$).
@@ VI Discussion and future work
In this article, we have presented a comprehensive overview of the ways in which Buchdahl’s limit can be bypassed. We have shown that the relaxation of one (or both) of the main assumptions of Buchdahl’s theorem—namely, the outward decreasing monotonicity of the density profile, and the isotropy of the stress-energy tensor—allows the existence of super-Buchdahl configurations. On the one hand, we have seen that, for non-decreasing (yet everywhere non-negative) density profiles, a new sort of Buchdahl limit seems to arise (as we illustrate for a toy model). We provide compelling evidence that the existence of a less compact bound than the one found in Baumgarte and Rendall (1993) is robust and independent of the model. We also show that as long as we allow negative densities (and hence violations of the weak energy condition), we can make objects as close to the compactness of a black hole as desired. Among the possibilities explored, we derived a particularly simple configuration consisting of an anti-de Sitter interior and a thin shell matching with an exterior Schwarzschild metric, that we have called AdS star, and which bares close resemblance with solutions obtained in semiclassical gravity Arrechea et al. (2024). On the other hand, we also introduced a very simple toy model of an anisotropic shell that allowed us to show that, without further restrictions (such as, e.g., energy conditions), one can conceive objects with compactness arbitrarily close to the black hole limit. The imposition of some energy condition still allows the surpassing of Buchdahl’s limit, but places new upper bounds on the compactness.
@@
When we allow non-monotonically decreasing density profiles, we may expect that if the matter content is that of an ordinary fluid, it will display an unstable behaviour, but no energy conditions need to be violated, a priori. If we go further and allow negative energy densities, causally pathological behaviour might arise due to the violation of the WEC. However, if such behaviour of the effective densities and pressures is motivated by an underlying healthy physical mechanism, e.g. semiclassical physics, the pathological behaviour might not arise. In fact, such behaviour is expected from assuming that perturbations with respect to the background constrain the density and pressures in a similar way the background ones do, without this being necessarily the case. For example, a horizonless alternative to GR black holes has recently appeared within semiclassical theories of gravity, where vacuum polarization effects in stars approaching Buchdahl’s limit predict the appearance of outwards increasing (and even negative) energy densities, allowing to bypass Buchdahl’s theorem Arrechea et al. (2024). The bilayered stars proposed in this work capture the core properties of these semiclassical solutions while being devoid of the intricacies of semiclassical analyses, which might foster their phenomenological scrutiny in the future.
@@ Acknowledgements.
The authors thank José M.M. Senovilla, Luis J. Garay, Ignacio Reyes and Ana Alonso-Serrano for helpful discussions. Financial support was provided by the Spanish Government through the Grants No. PID2020-118159GB-C43 and PID2023-149018NB-C43 (funded by MCIN/AEI/10.13039/501100011033), and by the Junta de Andalucía through the project FQM219. GGM is funded by the Spanish Government fellowship FPU20/01684. CB and GGM acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033. JPG acknowledges the support of a Mike and Ophelia Lazaridis Fellowship, as well as the support of a fellowship from “La Caixa” Foundation (ID 100010434, code LCF/BQ/AA20/11820043).
@@ Appendix B Fully anisotropic shell matching spherically symmetric spacetimes
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega_{2}.$ (64)
We can consider that the function $f(r)$ is piecewise defined, namely that we have:
$f(r)=\begin{cases}f_{-}(r)&r<R\\ f_{+}(r)&r>R\end{cases},$ (65)
where we will consider $f_{+}$ to be that of the Schwarzschild solution $f_{+}(r)=1-2M/r$, with $R\geq 2M$, and $f_{-}(r)$ to be the (A)dS and Minkowski solutions, namely $f_{-}(r)=1+\lambda r^{2}$, being AdS for $\lambda>0$, dS for $\lambda<0$, and Minkowski for $\lambda=0$. We can determine the distributional profile of the energy-momentum tensor that is required to match the two patches using Israel junction conditions.
@@
which translates into the condition $\sqrt{f_{+}(R)}t_{+}=\sqrt{f_{-}(R)}t_{-}$, choosing the same origin for the time coordinate. The vector tangent to the surface $u=\partial_{\tau}$ can be expressed as:
@@
$n^{+}_{a}=\frac{1}{\sqrt{f_{+}(r)}}\delta^{r}_{\ a},$ $n^{-}_{a}=\frac{1}{\sqrt{f_{-}(r+)}}\delta^{r}_{\ a}.$ (71)
@@
$K^{+}_{\tau\tau}=-\frac{1}{2}\frac{f^{\prime}_{+}}{\sqrt{f_{+}}},$ (75)
$K^{+}_{\theta\theta}=\sqrt{f_{+}}R,$ (76)
$K^{+}_{\Phi\Phi}=\sqrt{f_{+}}R\sin^{2}\theta,$ (77)
$K^{+}=K^{+}_{ab}h^{ab}=\frac{1}{2}\frac{f^{\prime}_{+}}{\sqrt{f_{+}}}+2\frac{\sqrt{f_{+}}}{R},$ (78)
whereas from the inside we simply need to replace $f_{+}$ by $f_{-}$:
$K^{-}_{\tau\tau}=-\frac{1}{2}\frac{f^{\prime}_{-}}{\sqrt{f_{-}}},$ (79)
$K^{-}_{\theta\theta}=\sqrt{f_{-}}R,$ (80)
$K^{-}_{\Phi\Phi}=\sqrt{f_{-}}R\sin^{2}\theta,$ (81)
$K^{-}=K^{-}_{ab}h^{ab}=\frac{1}{2}\frac{f^{\prime}_{-}}{\sqrt{f_{-}}}+2\frac{\sqrt{f_{-}}}{R}.$ (82)
@@
$\left[\left[K^{\tau}_{\ \tau}\right]\right]=\frac{1}{2}\frac{f^{\prime}_{+}}{\sqrt{f_{+}}}-\frac{1}{2}\frac{f^{\prime}_{-}}{\sqrt{f_{-}}},$ (84)
$\left[\left[K^{\theta}_{\ \theta}\right]\right]=\frac{\sqrt{f_{+}}}{R}-\frac{\sqrt{f_{-}}}{R},$ (85)
$\left[\left[K^{\Phi}_{\ \Phi}\right]\right]=\frac{\sqrt{f_{+}}}{R}-\frac{\sqrt{f_{-}}}{R},$ (86)
$\left[\left[K\right]\right]=\frac{1}{2}\left(\frac{f^{\prime}_{+}}{\sqrt{f_{+}}}-\frac{f^{\prime}_{-}}{\sqrt{f_{-}}}\right)+2\frac{\sqrt{f_{+}}-\sqrt{f_{-}}}{R}.$ (87)
@@
where $\sigma$ and $\tilde{p}_{t}$ represent the surface energy density and the surface tangential pressure, respectively, and are given by
$\sigma$ $=\frac{1}{4\pi R}\left(\sqrt{f_{-}(R)}-\sqrt{f_{+}(R)}\right),$ (89)
$\tilde{p}_{t}$ $=\frac{\sqrt{f_{+}(R)}-\sqrt{f_{-}(R)}}{8\pi R}+\frac{1}{16\pi R}\left(\frac{f^{\prime}_{+}}{\sqrt{f_{+}}}-\frac{f^{\prime}_{-}}{\sqrt{f_{-}}}\right).$ (90)

### v2

@@ IV AdS Stars and Einstein Static stars
AdS stars are constituted by an anti-de Sitter interior glued through a thin shell to a Schwarzschild exterior. The geometry can be built through a cut and paste procedure with the help of Israel junction conditions. The interesting thing here is that the qualitative behavior of this ad hoc geometry constructed through this surgery procedure appears naturally within the set of exact regular bilayered models, the only difference being that in that case the shell-like behavior is absent and the transition between the inside core and the outside is smooth.
@@
Let us specify the construction in detail: Take a spherical shell with radius $R$, such that its proper area is $4\pi R^{2}$. Let the spacetime outside the shell be Schwarzschild with a mass $M$, and AdS inside the shell. We denote the Schwarzschild (outside) patch with a $+$ sign, and the AdS (inside) patch with a -- sign. The metric is now given by the two line elements in Eq. (48)) with $f_{+}(r_{+})=1-2M/r_{+}$, $h_{+}(r_{+})=1/f_{+}(r_{+})$ and $f_{-}(r_{-})=1+\absolutevalue{\lambda}r^{2}$, $h_{-}(r_{-})=1/f_{-}(r_{-})$. The real parameter $\lambda$ corresponds to a negative cosmological constant. We locate the shell at $r_{+}=r_{-}=R$. The matching is worked out in detail in Appendix B, and we find that the shell displays a distributional stress-energy tensor with the following densities and pressures:
@@
These models could serve as simple templates of black hole mimickers inspired by semiclassical effects. In fact, more sophisticated models can be constructed with a nesting of two of these toy models, as done in Jampolski and Rezzolla (2024) with gravastars. One central element responsible for both the features of their ringdown Cardoso and Pani (2019) and their associated shadows Ayzenberg et al. (2023) are the redshift and time delay effects suffered by light rays crossing their interiors. While the AdS star has a redshift function that decreases towards the centre, the one associated to the Einstein static star is fully constant. This would result in radically different crossing times for null rays and, in turn, to very different observational signatures both in the electromagnetic and gravitational-wave spectra.
@@ V Anisotropy
For the sake of completeness in our analysis, in this section we will discuss the isolated effect of anisotropy in bypassing the Buchdahl limit. To this end, it is convenient to focus on a situation in which there are no violations of energy conditions arising from an interior core. For that purpose, we can analyze a toy model of a shell matching an exterior Schwarzschild spacetime with mass $M$ to an interior flat spacetime (i.e., Schwarzschild with $M=0$). The setup is the same as the one depicted in Fig. 10, but considering the core to be empty instead of AdS. This simply corresponds to setting the AdS constant $\lambda=0$ from previous results, and we find the following energy and tangential pressure densities:
@@
Given that the two of them have the same distributional behavior, for our particular example this would imply
@@
We can compare these bounds with the shell model that we are considering, and we realize that the shell model begins to violate the DEC whenever the compactness is greater than $0.962M/R=$ 0.96, which still obeys the bound $2M/R\leq 0.974$ found by Ivanov Ivanov (2002). Further results that put bounds on the compactness of the objects with similar assumptions are the results reported in Barraco et al. (2003); Boehmer and Harko (2006); Andréasson (2008); Urbano and Veermäe (2019), see also Rosa and Piçarra (2020). Similar analyses for elastic self-gravitating bodies were carried on in Alho et al. (2022a, b, 2024a, 2024b), reaching similar conclusions. In Barraco et al. (2003), under roughly speaking the same assumptions as Ivanov’s, it is shown that anisotropic models fall into two classes according to whether $p_{r}>p_{t}$ or $p_{t}>p_{r}$. In the case $p_{r}>p_{t}$, they manage to show that the $p_{r}$ function is strictly greater than the one in a fiduciary isotropic model with the same function $m(r)$. This means that these stars cannot be super-Buchdahl; actually, isotropic stars obeying DEC need to be less compact than Buchdahl Barraco and Hamity (2002). Regarding the case $p_{r}<p_{t}$, they show that the radial pressure is always bigger than the one in the fiduciary isotropic model with the same Misner-Sharp mass profile. Imposing the DEC, they derive bounds that are consistent with Ivanov’s. In Boehmer and Harko (2006), under some additional assumptions (about surface density behavior motivated on physical grounds), slightly sharper bounds are proved. Finally, in Lindblom (1984); Urbano and Veermäe (2019) it is shown that imposing a causality condition, namely that the speed of sound is smaller than the speed of light, there is a maximum compactness that can be reached, which turns out to be below Buchdahl’s limit, namely $2M/R\lesssim 0.768$.
In any case, the simple example exhibited in this section illustrates that in the absence of energy conditions, anisotropic configurations can be made as compact as desired. Energy conditions set a limit to the compactness, which is nevertheless far from Buchdahl’s, leaving the door open to having extremely compact objects that are close to forming event horizons, and are even allowed to display photon rings (since that only requires $2M/R>2/3\sim 0.666$).
@@ VI Discussion and future work
In this article, we have revisited the different ways in which Buchdahl’s limit can be bypassed, relaxing separately each of the main assumptions invoked in the theorem: namely, the outward-decreasing monotonicity of the density profile, and the isotropy of the stress-energy tensor. Regarding the former, we reviewed the bound obtained by Bondi Bondi (1964), which was later rigorously proven by Karageorgis and Stalker Karageorgis and Stalker (2008), showing that for non-decreasing (yet everywhere non-negative) density profiles, a compactness limit different from Buchdahl’s arises.
Here, we have found geometrically smooth toy models that allow us to approach this limit as closely as desired. We also show that as long as we allow negative densities (and hence violations of the weak energy condition), our toy model can describe objects arbitrarily close to the compactness of a black hole, in agreement with Bondi’s results. Among the possibilities explored, we derived a particularly simple configuration consisting of an anti-de Sitter interior and a thin shell matched to an exterior Schwarzschild metric, which we have called AdS star, and which bears a close resemblance to solutions obtained in semiclassical gravity Arrechea et al. (2024). On the other hand, we also introduced a very simple toy model of an anisotropic shell that allowed us to show that, without further restrictions (such as, e.g., energy conditions), one can conceive objects with compactness arbitrarily close to the black hole limit. The imposition of some energy condition still allows the surpassing of Buchdahl’s limit, but places new upper bounds on the compactness.
@@
When we allow non-monotonically decreasing density profiles, we may expect that if the matter content is that of an ordinary fluid, it will display an unstable behavior, but no energy conditions need to be violated, a priori. If we go further and allow negative energy densities, causally pathological behavior might arise due to the violation of the WEC. However, if such behavior of the effective densities and pressures is motivated by an underlying healthy physical mechanism, e.g. semiclassical physics, the pathological behavior might not arise. In fact, such behavior is expected from assuming that perturbations with respect to the background constrain the density and pressures in a similar way the background ones do, without this being necessarily the case. For example, a horizonless alternative to GR black holes has recently appeared within semiclassical theories of gravity, where vacuum polarization effects in stars approaching Buchdahl’s limit predict the appearance of outwards increasing (and even negative) energy densities, allowing to bypass Buchdahl’s theorem Arrechea et al. (2024). The bilayered stars proposed in this work capture the core properties of these semiclassical solutions while being devoid of the intricacies of semiclassical analyses, which might foster their phenomenological scrutiny in the future.
@@ Acknowledgements.
The authors thank José M.M. Senovilla, Luis J. Garay, Ignacio Reyes and Ana Alonso-Serrano for helpful discussions. We also thanks Shahar Hod and Håkan Andréasson for informing us about important references. Financial support was provided by the Spanish Government through the Grants No. PID2020-118159GB-C43 and PID2023-149018NB-C43 (funded by MCIN/AEI/10.13039/501100011033), and by the Junta de Andalucía through the project FQM219. GGM is funded by the Spanish Government fellowship FPU20/01684. CB and GGM acknowledge financial support from the Severo Ochoa grant CEX2021-001131-S funded by MCIN/AEI/ 10.13039/501100011033. JPG acknowledges the support of a Mike and Ophelia Lazaridis Fellowship, as well as the support of a fellowship from “La Caixa” Foundation (ID 100010434, code LCF/BQ/AA20/11820043).
@@ Appendix B Fully anisotropic shell matching spherically symmetric spacetimes
$ds^{2}=-f(r)dt^{2}+h(r)dr^{2}+r^{2}d\Omega_{2}.$ (84)
We can consider that the function $f(r)$ and $h(r)$ are piecewise defined, namely that we have:
$f(r)=\begin{cases}f_{-}(r)&r<R\\ f_{+}(r)&r>R\end{cases},\qquad h(r)=\begin{cases}h_{-}(r)&r<R\\ h_{+}(r)&r>R\end{cases},$ (85)
@@
which translates into the condition $\sqrt{f_{+}(R)}t_{+}=\sqrt{f_{-}(R)}t_{-}$, choosing the same origin for the time coordinate. In other words, the induced metric is given by
$ds^{2}=-d\tau^{2}+R^{2}d\Omega^{2},$ (89)
since we can identify the angular coordinates and the proper time long the shell is adapted to the redshift functions from the inside and the outside. The vector tangent to the surface $u=\partial_{\tau}$ can be expressed as:
@@
$n^{+}_{a}=\sqrt{h_{+}(r)}\delta^{r}_{\ a},$ $n^{-}_{a}=\sqrt{h_{-}(r)}\delta^{r}_{\ a}.$ (92)
@@
$K^{+}_{\tau\tau}=-\frac{1}{2}\frac{f^{\prime}_{+}(R)}{f_{+}(R)\sqrt{h_{+}(R)}},$ (96)
$K^{+}_{\theta\theta}=\frac{R}{\sqrt{h_{+}(R)}},$ (97)
$K^{+}_{\Phi\Phi}=\frac{R\sin^{2}\theta}{\sqrt{h_{+}(R)}},$ (98)
$K^{+}=K^{+}_{ab}h^{ab}=\frac{1}{2}\frac{f^{\prime}_{+}(R)}{f_{+}(R)\sqrt{h_{+}(R)}}+2\frac{R}{\sqrt{h_{+}(R)}},$ (99)
whereas from the inside we simply need to replace $f_{+}$ and $h_{+}$ by $f_{-}$ and $h_{-}$ respectively:
$K^{-}_{\tau\tau}=-\frac{1}{2}\frac{f^{\prime}_{-}(R)}{f_{-}(R)\sqrt{h_{-}(R)}},$ (100)
$K^{-}_{\theta\theta}=\frac{R}{\sqrt{h_{-}(R)}},$ (101)
$K^{-}_{\Phi\Phi}=\frac{R\sin^{2}\theta}{\sqrt{h_{-}(R)}},$ (102)
$K^{-}=K^{-}_{ab}h^{ab}=\frac{1}{2}\frac{f^{\prime}_{-}(R)}{f_{-}(R)\sqrt{h_{-}(R)}}+2\frac{R}{\sqrt{h_{-}(R)}},$ (103)
@@
$\left[\left[K^{\tau}_{\ \tau}\right]\right]=\frac{1}{2}\frac{f^{\prime}_{+}(R)}{f_{+}(R)\sqrt{h_{+}(R)}}-\frac{1}{2}\frac{f^{\prime}_{-}(R)}{f_{-}(R)\sqrt{h_{-}(R)}},$ (105)
$\left[\left[K^{\theta}_{\ \theta}\right]\right]=\frac{1}{R\sqrt{h_{+}(R)}}-\frac{1}{R\sqrt{h_{-}(R)}},$ (106)
$\left[\left[K^{\Phi}_{\ \Phi}\right]\right]=\frac{1}{R\sqrt{h_{+}(R)}}-\frac{1}{R\sqrt{h_{-}(R)}},$ (107)
$\left[\left[K\right]\right]=\frac{1}{2}\left(\frac{f^{\prime}_{+}(R)}{f_{+}(R)\sqrt{h_{+}(R)}}-\frac{f^{\prime}_{-}(R)}{f_{-}(R)\sqrt{h_{-}(R)}}\right)+\frac{2}{R}\left[\frac{1}{\sqrt{h_{+}(R)}}-\frac{1}{\sqrt{h_{-}(R)}}\right].$ (108)
@@
where $\sigma$ and $\tilde{p}_{t}$ represent the surface energy density and the surface tangential pressure. We have $S^{\tau}_{\ \tau}=-\sigma$ and $S^{\theta}_{\ \theta}=\tilde{p}_{t}$. They are given by
$\sigma$ $=\frac{1}{4\pi R}\left(\frac{1}{\sqrt{h_{-}(R)}}-\frac{1}{\sqrt{h_{+}(R)}}\right),$ (110)
$\tilde{p}_{t}$ $=\frac{1}{8\pi R}\left[\frac{1}{\sqrt{h_{+}(R)}}-\frac{1}{\sqrt{h_{-}(R)}}\right]+\frac{1}{16\pi}\left(\frac{f^{\prime}_{+}(R)}{f_{+}(R)\sqrt{h_{+}(R)}}-\frac{f^{\prime}_{-}(R)}{f_{-}(R)\sqrt{h_{-}(R)}}\right).$ (111)
In general, whenever $\tilde{p}_{t}\neq 0$, the presence of a thin-shell automatically introduces anisotropic pressures. However, the special situations in which $\tilde{p}_{t}=0$ do not introduce any anisotropic pressure into play, since the only component of the energy-momentum tensor that does not identically vanish corresponds to the energy density, which exhibits a distributional behavior.